# Perverse microsheaves

### v2

@@ 7. The perverse t-structure
The prototypical example is when $\mathcal{T}$ is a derived category of chain complexes, and $\mathcal{T}^{\leq 0}$ consists of the complexes whose cohomology is concentrated in degrees $\leq 0$, etc.
@@
The notion of $t$-structure is imported to the setting of stable categories in [19, Sec. 1.2]: By definition, a $t$-structure on a stable category is a $t$-structure on its homotopy category, which canonically carries the structure of a triangulated category. It is shown that the various properties of $t$-structures lift to the stable setting, in particular, the existence of truncation functors, and the fact that the full subcategory on objects in the heart is abelian. [^4] Let us avoid a possible source of confusion. One might think that, insofar as stable categories generalize dg categories, the heart could be expected to have, in its hom spaces, whatever corresponds to the positive ext groups. This depends on whether or not the stable category is viewed as a usual $\infty$-category, or as an $\infty$-category enriched in spectra. Indeed, the positive ext groups (in cohomological grading conventions) correspond to negative homotopy groups, so are only manifest after the (canonical) enrichment in spectra. Here however the statement about the heart should be understood in terms of the not enriched $\infty$-category. For consistency with [3] (and in contrast to [19]), we adopt cohomological conventions and write $H^{i}$ instead of $\pi_{-i}$.
@@
When $C$ is a presentable stable category, then if either $C^{\geq 0}$ or $C^{\leq 0}$ is presentable, then so is the other, and all truncation functors are colimit preserving [19, 1.4.4.13]. In this case, the subcategory of compact objects $C^{c}$ is stable under the truncation functors and inherits a $t$-structure. Indeed, $\tau^{\geq 0}$ is left adjoint to the corresponding inclusion, assumed colimit preserving, hence $\tau^{\geq 0}$ preserves compact objects. Taking cones, so does $\tau^{\leq 0}$.
@@
It is easy to see that the property that $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure may be checked on sections on any base of open sets. Indeed, regarding condition (i) and (ii) of Definition 7.1, it is immediate from the sheaf condition that vanishing of Homs and containment of subcategories can be checked locally.
Regarding (iii), the key point is that for any candidate $t$-structure satisfying (i) and (ii), the space of fiber sequences $K^{\prime}\to K\to K^{\prime\prime}$ as requested in (iii) is either empty or contractible. Indeed, first recall that using property (ii) to apply (i) to shifts, we find that the following strengthening of (i): for any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, the negative exts, aka positive homotopy groups of the hom space $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])$, must vanish. Now given any $K^{\prime}\to K\to K^{\prime\prime}[-1]$ and $L^{\prime}\to K\to L^{\prime\prime}[-1]$ both satisfying (iii), we obtain a canonical null-homotopy of the composition $K^{\prime}\to K\to L^{\prime\prime}[-1]$ hence lift of $K^{\prime}\to K$ to $K^{\prime}\to L^{\prime}$, etc.
Having learned this contractibility, if (iii) holds locally, then we can canonically glue the local exact triangles to obtain (iii) globally.
@@
We recall a comparison between various notions of sheaves of categories. For covers $U=\bigcup U_{i}$, an $\infty$-categorical sheaf of categories $\mathcal{C}$ satisfies the condition that
@@
$\mathcal{C}(U)=\lim\left(\prod_{i\in I}\mathcal{C}(U_{i})\rightrightarrows\prod_{i,j\in I}\mathcal{C}(U_{i}\cap U_{j})\mathrel{\begin{subarray}{c}\textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\end{subarray}}\prod_{i,j,k\in I}\mathcal{C}(U_{i}\cap U_{j}\cap U_{k})\cdots\right)$
The corresponding definition of stacks in the classical literature instead involves a limit of 1-categories taken in the $(2,1)$-category of ordinary categories. (Note that for limits in $(2,1)$ categories, the compatibility condition on triple overlaps is already strict, hence the terms after triple overlaps in the above diagram are irrelevant to the limit. In the literature they are correspondingly not included in the descent condition.
The situation is analogous to the fact that the descent condition for sheaves of sets is usually given truncated after the second term.)
These notions agree: $1$-categories are $1$-truncated objects of $(\infty,1)$-categories, and, for any category $\mathcal{C}$, the inclusion of $k$-truncated objects $\tau_{\leq k}\mathcal{C}\hookrightarrow\mathcal{C}$ is limit-preserving [18, Proposition 5.5.6.5]. We summarize:
@@
The heart of a $t$-structure on a sheaf of stable categories is a stack of abelian categories.
7.2. The perverse t-structure on constructible sheaves
We now review from [15, Sec. 10.3] the microlocal description of the perverse $t$-structure on constructible sheaves.
Let $M$ be a complex manifold. For a Lagrangian subset $\Lambda\subset T^{*}M$, we write $\Lambda^{cl}$ for the locus of smooth points of $\Lambda$ where the map $\Lambda\to M$ has locally constant rank. For a point $p\in\Lambda^{cl}$, we write $r(p)$ for this rank. Fix a $t$-structure $C^{\leq 0},C^{\geq 0}$ on our coefficient category $C$.
We write [15, Def. 10.3.7] using the notation of (10):
@@ 7.2. The perverse t-structure on constructible sheaves
[15, Def. 10.3.7] Consider the following full subcategories of $sh(M)_{\mathbb{C}-c}$.
${}^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)]\in C^{\leq 0}\}$
${}^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)])\in C^{\geq 0}\}$
It is proved in [15, Thm. 10.3.12] that the above pair agrees with the usual definition of the perverse $t$-structure on $M$, and so in particular is a $t$-structure. [^5] The result is stated in [15] for $C$ the bounded derived category of modules over a ring and and $t$ the standard $t$-structure. However, the arguments given there (or in [3]) for the existence of the perverse $t$-structure depend only on the general properties of the six functor formalism, and the comparison between Definition 7.7 and the usual stalk/costalk wise definition of the perverse $t$-structure depends only on standard properties of microsupports.
@@
If $\Lambda\subset T^{*}M$ is (possibly singular) subanalytic complex Lagrangian then Definition 7.7 induces a $t$-structure on $sh_{\Lambda}(M)$. Moreover, $sh_{\Lambda}(M)^{\heartsuit}$ is closed under extensions inside $sh(M)$.

### v3

@@ 6. The perverse t-structure
The prototypical example is when $\mathcal{T}$ is a derived category of chain complexes, $\mathcal{T}^{\leq 0}$ (resp. $\mathcal{T}^{\geq 0}$) consists of the complexes whose cohomology is concentrated in degrees $\leq 0$ (resp. $\geq 0$).
@@
The notion of $t$-structure is imported to the setting of stable categories in [30, Sec. 1.2]: By definition, a $t$-structure on a stable category is a $t$-structure on its homotopy category, which canonically carries the structure of a triangulated category. It is shown that the various properties of $t$-structures lift to the stable setting, in particular, the existence of truncation functors, and the fact that the full subcategory on objects in the heart is abelian. [^3] Let us avoid a possible source of confusion. One might think that, insofar as stable categories generalize dg categories, the heart could be expected to have, in its hom spaces, whatever corresponds to the positive ext groups. This depends on whether or not the stable category is viewed as a usual $\infty$-category, or as an $\infty$-category enriched in spectra. Indeed, the positive ext groups (in cohomological grading conventions) correspond to negative homotopy groups, so are only manifest after the (canonical) enrichment in spectra. Here however the statement about the heart should be understood in terms of the not enriched $\infty$-category. For consistency with [3] (and in contrast to [30]), we adopt cohomological conventions and write $H^{i}$ instead of $\pi_{-i}$.
@@
When $\mathcal{C}$ is a presentable stable category, then if either $\mathcal{C}^{\geq 0}$ or $\mathcal{C}^{\leq 0}$ is presentable, then so is the other, and all truncation functors are colimit preserving [30, 1.4.4.13]. In this case, the subcategory of compact objects $\mathcal{C}^{c}$ is stable under the truncation functors and inherits a $t$-structure. Indeed, $\tau^{\geq 0}$ is left adjoint to the corresponding inclusion, assumed colimit preserving, hence $\tau^{\geq 0}$ preserves compact objects. Taking cones, so does $\tau^{\leq 0}$.
@@
Lemma 6.5.
The property that $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure may be checked on sections on any base of open sets.
Proof.
Indeed, regarding condition (i) and (ii) of Definition 6.1, it is immediate from the sheaf condition that vanishing of Homs and containment of subcategories can be checked locally.
Regarding (iii), the key point is that for any candidate $t$-structure satisfying (i) and (ii), the space of fiber sequences $K^{\prime}\to K\to K^{\prime\prime}$ as requested in (iii) is either empty or contractible. Indeed, first recall that using property (ii) to apply (i) to shifts, we find the following strengthening of (i): for any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, the negative exts, aka positive homotopy groups of the hom space $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])$, must vanish. Now given any $K^{\prime}\to K\to K^{\prime\prime}[-1]$ and $L^{\prime}\to K\to L^{\prime\prime}[-1]$ both satisfying (iii), we obtain a canonical null-homotopy of the composition $K^{\prime}\to K\to L^{\prime\prime}[-1]$ hence lift of $K^{\prime}\to K$ to $K^{\prime}\to L^{\prime}$, etc.
Having learned this contractibility, if (iii) holds locally, then we can canonically glue the local exact triangles to obtain (iii) globally. ∎
@@
In the classical literature, the sheaf condition for sheaves of 1-categories (such sheaves are sometimes called stacks) is formulated as a limit of 1-categories taken in the $(2,1)$-category of ordinary categories. In this formulation, the compatibility condition on triple overlaps is strict.
@@
By contrast, the notion of $\infty$-categorical sheaf of $(\infty,1)$-categories requires that for such a sheaf $\mathcal{C}$ and covers $U=\bigcup U_{i}$, one has
@@
$\mathcal{C}(U)=\lim\left(\prod_{i\in I}\mathcal{C}(U_{i})\rightrightarrows\prod_{i,j\in I}\mathcal{C}(U_{i}\cap U_{j})\mathrel{\begin{subarray}{c}\textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\end{subarray}}\prod_{i,j,k\in I}\mathcal{C}(U_{i}\cap U_{j}\cap U_{k})\cdots\right).$
These notions are equivalent for $1$-categories: one passes from the $\infty$-categorical notion to the $1$-categorical notion by truncation, and for the reverse direction one need only note that $1$-categories are $1$-truncated objects of $(\infty,1)$-categories, and the inclusion of $k$-truncated objects is limit-preserving [32, Proposition 5.5.6.5].
6.2. The perverse t-structure on constructible sheaves
We now review from [25, Sec. 10.3] the microlocal description of the perverse $t$-structure on constructible sheaves.
Let $M$ be a complex manifold. For a Lagrangian subset $\Lambda\subset T^{*}M$, we write $\Lambda^{\circ}$ for the locus of smooth points of $\Lambda$ where the map $\Lambda\to M$ has locally constant rank. Fix a $t$-structure $\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}$ on our coefficient category $\mathcal{C}$, with corresponding truncation functors $\tau^{\leq 0},\tau^{\geq 0}$. It is proved in [25, Theorem 10.3.12] that the following prescription characterizes the perverse $t$-structure on $sh(M)_{\mathbb{C}-c}$. [^4] The results stated in [25] are for $\mathcal{C}$ the bounded derived category of modules over a ring and and $t$ the standard $t$-structure. However, the arguments given there (or in [3]) for the existence of the perverse $t$-structure depend only on the general properties of the six functor formalism, and the comparison between Definition 6.7 and the usual stalk/costalk wise definition of the perverse $t$-structure depends only on standard properties of microsupports.
Definition 6.7 ([25, (10.3.7) and Definition 10.3.7]).
Let $^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}$ (resp. $^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}$) be the full subcategory of $sh(M)_{\mathbb{C}-c}$ on objects $F$ with the property that, for every $p\in ss(F)^{\circ}$ such that $\pi:SS(F)\to M$ has constant rank on a neighborhood of $p$, there exists a submanifold $N$ and $L\in\mathcal{C}$ such that $F\simeq L_{Y}[dim(N)]\in(\mu sh_{T^{*}M})_{p}$ and $\tau^{\geq 1}F\simeq 0$ (resp. $\tau^{\leq-1}F\simeq 0$).
Definition 6.7 can be equivalently expressed using the microstalk functor of Corollary 1.3.
Definition 6.8.
Consider the following full subcategories of $sh(M)_{\mathbb{C}-c}$.
$^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{\circ}\implies\omega_{p}^{-1}F[-n]\in\mathcal{C}^{\leq 0}\}$
$^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{\circ}\implies\omega_{p}^{-1}F[-n]\in\mathcal{C}^{\geq 0}\}$
Proposition 6.9.
Proof.
Follows immediately from Corollary 5.11. ∎
@@ 6.2. The perverse t-structure on constructible sheaves
If $\Lambda\subset T^{*}M$ is (possibly singular) subanalytic complex Lagrangian and $\Omega\subseteq\mathbb{P}^{*}M$ is an open set, then Definition 6.7 induces a $t$-structure on
$sh_{\Lambda\cup\pi^{-1}(\Omega^{c}),\mathbb{C}-c}(M)=\{F\in sh_{\mathbb{C}-c}(M)|ss(F)\cap\pi^{-1}(\Omega)\subseteq\Lambda\}.$
Moreover, $sh_{\Lambda\cup\pi^{-1}(\Omega^{c}),\mathbb{C}-c}(M)^{\heartsuit}$ is closed under extensions inside $sh(M)$.
Proof.
By (1) of Lemma 6.2, we only need to check that $\tau_{\leq k}F$ is contained in the subcategory if $F$ is. Suppose not; then for any neighborhood $U$ of $\mathbb{C}^{*}\cdot p$, $ss(\tau_{\leq k}F)-\Lambda$ is non-empty. Since $\tau_{\leq k}F$ is constructible, $(ss(\tau_{\leq k}F)-\Lambda)\cap U$ must have a smooth Legendrian point $q$. But then the microstalk of $\tau_{\leq k}F$ at this point $q$ is the truncation of the microstalk of $F$, which is zero. A contradiction.
Theorem 6.11.
[42] The microlocal perverse t-structure in Definition 6.8 induces a perverse $t$-structure on $\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$. Furthermore, for $\mathcal{F},\mathcal{G}\in[\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})]^{\heartsuit}$, the hom-sheaf
$\mathcal{H}om_{\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})}(\mathcal{G},\mathcal{F})[\dim M]$
is a perverse sheaf on the $T^{\circ}M$.
Proof.
By Proposition 5.9, the inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}\xhookrightarrow{\sim}\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$ is an equivalence so we can work with $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}$. Any $\mathcal{F}\in\operatorname{\mathbb{P}\mu sh}_{P^{*}M,\mathbb{C}-c}(\Omega)$ tautologically belongs to $\operatorname{\mathbb{P}\mu sh}_{\Lambda}(\Omega)$ where $\Lambda\coloneqq\operatorname{supp}(\mathcal{F})$, so we could fix a support condition $\Lambda$. By Lemma 6.5, it’s enough to check it on an open cover. But since $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$ is constructible, there exits a cover $\mathcal{U}$ such that each $\Omega$ satisfies $\operatorname{\mathbb{P}\mu sh}_{\Lambda}(\Omega)=(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}$ for some $p\in\Omega$, so we can check on stalks. By Corollary 5.8 and Proposition 6.9, but $(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}$ and the notion of microstalks are invariant under contact transform, so we can assume $\Lambda$ is in general position. In this case, by Theorem 5.3, $(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}\hookrightarrow(sh_{\Lambda})_{m}/loc_{m}$ has image given by the category in the previous Lemma 6.10, which we’ve to have a t-structure, and it clear induces a $t$-structure on the quotient by applying Lemma 6.2 to locloc.
For the statement regarding the hom sheaf, let $\mathcal{F},\mathcal{G}\in(\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c})^{\heartsuit}$. Since perversity can be checked locally, we may pick some sheaves $F$ and $G$ representing $\mathcal{F}$ and $\mathcal{G}$. But in this case, by (10), we have the identification
$\mathcal{H}om_{\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c}}(\mathcal{G},\mathcal{F})=\mathcal{H}om_{\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c}}(G,F)=\mu hom(G,F).$
But then, [25, Corollary 10.3.20] implies that $\mathcal{H}om_{\mathbb{P}\mu sh_{V}}(\mathcal{G},\mathcal{F})[\dim M]=\mu hom(G,F)[\dim M]$ is perverse. ∎
We now define notion of perverse t-structure for microsheaves on complex symplectic and contact manifolds for the canonical microsheaves over $\mathcal{C}=R-mod$, for a discrete commutative ring $R$. We postpone the general discussion later in Definition 6.16 for conceptual clarity.
Recall by Corollary 1.3, for any microsheaf $\mathcal{F}$ on $V$, there is a well-defined object $\omega_{p}^{-1}(F)\in\mathcal{C}$.
@@ 6.3. Perverse microsheaves on complex contact and symplectic manifolds
Let $V$ be a contact manifold of complex dimension 2n12n-1. We define the pair of subcategories $((\mu sh_{\widetilde{V},\mathbb{C}-c})^{\leq 0},(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\geq 0})$ by constraining the microstalks:
$(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\leq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\leq 0}\}$
$(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\geq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\geq 0}\}$
We define similarly the corresponding notions for objects supported in some fixed (singular) Lagrangian, and define as always the corresponding notions on $V$ by pushforward.  [^5] The $[n]$ is just a convention, set to match the usual conventions for perverse sheaves.
Theorem 6.13.
$(\mu sh_{V,\mathbb{C}-c})^{\leq 0}$, $(\mu sh_{V,\mathbb{C}-c})^{\geq 0}$ determine a $t$-structure on $\mu sh_{V,\mathbb{C}-c}$. In particular, $(\mu sh_{V,\mathbb{C}-c})^{\heartsuit}$ is a sheaf of abelian categories.
Furthermore, for $\mathcal{F},\mathcal{G}\in\mu sh_{V,\mathbb{C}-c}^{\heartsuit}$, the sheaf of morphisms
$\mathcal{H}om_{\mu sh_{V}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim\widetilde{V}$]
is a perverse sheaf on the symplectization $\widetilde{V}$.
Proof.
Per Lemma 6.5, a pair of subcategories $(\mathcal{C}^{\leq},\mathcal{C}^{\geq})$ being a t-structure can be checked on open covers. Similarly, being a perverse sheaf if also a local condition. Thus, we reduce to Theorem 6.11 by taking Darboux charts. ∎
We deduce Theorem 1.6 from the introduction.
Proof of Theorem 1.6.
By Theorem 5.16, there is an equivalence $(({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}})^{\mathbb{C}^{*}}\simeq({\gamma_{\mathbb{C}}})_{*}\mu sh_{W}.$ The $\mathbb{C}^{*}$-action manifestly preserves the subspaces $\mu sh_{W,\mathbb{C}-c}^{\leq 0}$ and $\mu sh_{W,\mathbb{C}-c}^{\geq 0}$, so the $t$-structure passes to the invariants (taking hearts is a pullback ($\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$) hence a limit, and hence commutes with taking $G$-invariants which is also a limit).
Now, we consider the perversity of sheaf Hom. We recall that objects here admits two different interpretations: as microsheaves on $\widetilde{{W\times\mathbb{C}}}$ and as microsheaves on the underlying real $W$ since Theorem 5.16 is needed to define $\mu sh_{W,\mathbb{C}-c}^{\heartsuit}$. Denote by $\widetilde{\mathcal{F}}$, $\widetilde{\mathcal{G}}$ for the former and $\mathcal{F}$, $\mathcal{G}$ for the latter, then we see from the contact case that $\mu hom(\widetilde{\mathcal{G}},\widetilde{\mathcal{F}})[\frac{1}{2}(\dim W)+1)]$ is a perverse sheaf on $\widetilde{W\times\mathbb{C}}$. However, to descend them onto $W$, we recall that such objects are assumed to be locally constant along the $\mathbb{C}^{*}$-action and we take $\mathbb{C}^{*}$-invariant to quotient the extra direction. This process drops the dimension by $1$ and we thus conclude that $\mu hom(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim W]$ is perverse. ∎
Now, let $\mathcal{C}$ be any symmetric monoidal category and $(\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0})$ be a $t$-structure on $\mathcal{C}$. We explain how one can generalize Definition 6.12 to allow more flexibility on Maslov data. For this purpose, we recall the notion of constrained Maslov data from Section 4.7: A collection of subcategories (of $\mathcal{C}$) $\{\mathcal{D}_{i}\}$ is said to be anchored if the submonoid $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$ fixing each $\mathcal{D}_{i}$ is a subgroup.
@@
Let $(\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0})$ be a $t$-structure on $\mathcal{C}$, then the collection $\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\}$ is anchored.